Corect Part C The charges are in a uniform electric field whose direction makes an angle 36 1 with the line connecting the charges. What is the magnitude of this field if the torque exerted on the dipole has magnitude 720 10 N-m Express your answer in newtons per coulomb to three significant figures. View Available Hint(s) N/C
Answer #1

Part C )

given

angle \theta = 36.1o  

the torque is \tau = 7.2 x 10-9 N.m

but here the formula is using to find \tau is

\tau = p E sin\theta

E = \tau / p sin\theta

here p = q d

both values of q and d are not given so plug in q and d values and get ' p ' value then

substitute as following,

E = \tau / p sin\theta

= 7.2 x 10-9 / p x sin36.1

= 1.22 x 10-8 / p ( substitute value )

E = 1.22 x 10-8 / p   N/C